OpenStudy (anonymous):

Help Please. Need somebody to help explain how to simplify expressions. 9. 3^-2 I did this: 1 divided by ( 3 x 3) which i got 1/9 ....

OpenStudy (anonymous):

No like this. 

OpenStudy (jacobbenvenutty):

Just do 1 divided by 9

OpenStudy (anonymous):

its 0.11111111

OpenStudy (jacobbenvenutty):

Example: 5-2 52=1 * 5 * 5 = 25 5-2=1 / 25=0.04

OpenStudy (anonymous):

so would 0.1111111 be the simplified expression?

OpenStudy (jacobbenvenutty):

Yes

OpenStudy (anonymous):

Multiply the base number (3) by itself two times (because the exponent is a 2)

OpenStudy (anonymous):

Ohh Ohkay.

OpenStudy (jacobbenvenutty):

She did already : )

OpenStudy (anonymous):

So to make sure i got this . my next one is (-5)^-2 i did (-5 x -5 x -5) and got -125 so it would 1/-125 and it would -0.008 ??

OpenStudy (anonymous):

@Jacobbenvenutty

OpenStudy (jacobbenvenutty):

Since it is only to the second power and not the third power it would look like this: 1/(-5 x -5)= 1/25

OpenStudy (anonymous):

Oh that's what i meant!

OpenStudy (anonymous):

What about this one?